OpenStudy (anonymous):

How do you take the derivative of this?? f(x)=1/x^2

OpenStudy (bahrom7893):

rewriting it as f(x)=x^(-2) makes it a piece of cake.

OpenStudy (bahrom7893):

f'(x)=-2*x^(-3)

OpenStudy (bahrom7893):

or: f'(x)=-2/x^3

OpenStudy (kainui):

x^-1=1/x So any time you see something written like that, you can just flip it around and give it a negative exponent. If you do, you can avoid using the quotient rule!

OpenStudy (anonymous):

so how do u solve if f'(4)??

OpenStudy (anonymous):

Thanks i got it just plug in 4